# Weaver: Interweaving SQL and LLM for Table Reasoning

### v1

@@ 3.2 Results and Analysis
Table 4: Number of API calls comparison per TableQA.
ProTrix uses only two fixed API calls and relies on the LLM solely for generating the plan. However, it does not involve the LLM during execution. This limits its ability to handle multi-step queries requiring reasoning at each step. For instance, it may fail to infer information from individual rows or perform operations such as applying a SQL GROUP BY on an LLM-inferred ’country’ column. Such steps are often essential to arrive at the correct final answer. These metrics demonstrate how our approach minimizes computational overhead while maintaining accuracy.
@@
Optimization:
We experimented with optimizing our planning and execution strategy on 200 TableQA queries. Table 5 demonstrates how our optimization strategy focuses on reducing unnecessary computational steps without compromising accuracy. To achieve this, we implemented several techniques.
1. LLM Step Reduction: By identifying and eliminating redundant LLM steps, we reduced the 15 LLM calls. This optimization ensures that LLMs are only used when necessary, lowering computational costs.
2. SQL Step Optimization: We achieved a reduction of 19 SQL steps by eliminating unused operations. Additionally, we merged 113 sequential SQL steps into fewer, more efficient queries and reordered 4 steps to optimize execution flow, making it more efficient.
@@
Table 6: Error in SQL and Plan Generation on WikiTQ.
@@
We perform a stage-wise analysis to assess the contribution of each component in our pipeline—filtering, planning, and execution. Compared to SQL-only generation, which struggles with multi-step reasoning, our pipeline yields consistent accuracy gains by structuring the task into sub-components. The filtering stage removes irrelevant columns, reducing noise and guiding the model’s attention. The planning stage—central to our method—decomposes complex queries into symbolic and semantic steps. This step is essential and not ablatable. However, skipping plan verification (i.e., executing without validating) leads to a 1% drop in accuracy, indicating that verification adds robustness. Finally, execution stage translates structured plans into SQL.
4 Comparision with Related Work
Table-based Question Answering (TableQA) combines table understanding, question interpretation, and NLP. Foundational work such as Text-to-SQL Rajkumar et al. (2022), Program-of-Thought Chen et al. (2023), and TabSQLify Nahid and Rafiei (2024b) laid the groundwork. Binder and TAG Biswal et al. (2024) expose the limitations of traditional Text-to-SQL methods in handling complex analytical tasks involving both structured and unstructured data. To address these challenges, several alternative approaches have been explored:
Fine-Tuning Methods: These methods fine-tune LLMs to specialize in reasoning over hybrid tabular and textual data. Models such as Zhu et al. (2024), Mittal et al. (2024), and Patel et al. (2024) are trained to extract, reason, and execute over such inputs. However, fine-tuning requires large task-specific datasets and tends to lack generalization across domains.
Query Engines: This direction integrates LLMs with SQL engines via user-defined functions (UDFs), allowing LLM calls within queries. UQE Dai et al. (2024), BlendSQL, SUQL Liu et al. (2024a), and Binder follow this paradigm. While flexible, LLM-generated queries can be error-prone, and these systems often support only limited query structures, reducing adaptability.
Table Pruning and Planning: Approaches like H-Star, ReAcT, ProTrix, and others Liu et al. (2024b); Nahid and Rafiei (2024a) enhance efficiency by programmatically pruning rows or columns using SQL or Python. While this reduces processing overhead, these methods often function as black boxes, lacking transparency and vulnerable to cascading errors if early pruning steps are incorrect.
@@ 5 Conclusion
We introduce Weaver , a novel approach for table-based question answering on semi-structured tables. Weaver outperforms all baselines by strategically decomposing complex queries into a sequence of LLM- and SQL-based planning steps. By alternating between these modalities, it enables precise, interpretable, and adaptive query resolution. Weaver overcomes prior limitations by effectively handling both complex queries and large tables. Its modular design also supports future extensions, including image-based tables, multi-table reasoning, and integration with free-form text. As future work, we plan to explore fine-tuning and supervision strategies to further improve execution accuracy and plan reliability.
@@ Acknowledgments
We gratefully acknowledge the Cognitive Computation Group at the University of Pennsylvania and the Complex Data Analysis and Reasoning Lab at Arizona State University for their resources and computational support.
@@ 2. The list will be used to filter the table dataframe directly so take care of that
- Check if the MySQL step logic adheres to the column format. (Performs calculations and formatting and filtering in the table)
@@
Do not provide with code or other explanations, only the new plan.
@@
1. Store the output at each step by creating a new table. Use this new table for the next steps.
2. The code for MySQL should handle all values in the column (formatting and filtering). New columns
@@
LLM step format should be same.
@@
Give me code for solving the question, and no other explanations. Keep in mind the column data formats (string to appropriate data type, removing extra character, Null values) while writing Mysql code.
@@
- Do not provide any explanation and Return only a list (separate values by ’#’) that can be added
@@
- Any value should not be more than 3 words (or each value should be as short as possible).
@@
If the table has just one column or value consider that as the answer given the column name.
@@
Listing 8: Plan Optimization Prompt
@@
SQL Merge: Merge sequential SQL steps where possible (e.g., combining filtering, aggregation, and sorting in one query).
SQL Reordering: Reorder SQL steps to filter early before applying computationally expensive operations like LLM processing.
@@ A.2 Table and Question Example
Listing 9: Table and Question Example for LLM Planning
@@
Question: How long did it take for the New York Americans to win the National Cup after 1936?
@@ A.3 Model Responses
Listing 10: Response to Sample Prompt 1 (Column Description)
@@
The table New_York_Americans_soccer contains historical performance data for the New York Americans soccer team, detailing their standings in various leagues, playoff outcomes, and national cup results from 1931 to 1956. The data includes information on the year, division, league, regular season standings, playoff participation, and national cup results.
@@
| Year | String | Standardize to a consistent format | Represents the year or season of the soccer performance. Some entries may need parsing to extract the year. |
| Division | Float | Convert to Integer (if applicable) | Indicates the division in which the team played. Some entries are NaN and need handling. |
| League | String | Standardize to a consistent format | Represents the league in which the team participated, primarily ASL (American Soccer League). |
| Reg_Season | String | Clean and standardize standings | Indicates the teams regular season performance, which may include qualifiers like "t" or "?" that need clarification. |
| Playoffs | String | Standardize and clean | Indicates playoff participation and outcomes, with entries like "No playoff" and "Finals" that need to be standardized. |
| National_Cup | String | Standardize and clean | Indicates the outcome of the national cup, with entries like "Champion" or "1st Round" that need standardization. |
Listing 11: Response to Sample Prompt 2 (Planning Prompt)
Plan: Step 1: SQL - Standardize the Year column to a consistent format and extract the year from
entries like "Spring 1932" and "Fall 1932".
Step 2: SQL - Clean and standardize the National_Cup column to identify the years when the team won
the national cup.
Step 3: SQL - Filter the data to find the first year after 1936 when the National_Cup column
indicates a win.
Listing 12: Response to Sample Prompt 3 (Verify Plan Prompt)
@@
Step 1: LLM - Standardize the Year column to a consistent format by extracting the year from entries
like "Spring 1932" and "Fall 1932". Convert all entries to a four-digit year format (e.g., "1932"
instead of "Spring 1932").
Step 2: SQL - Clean and standardize the National_Cup column to identify winning entries. Define a
clear criterion for a "win," such as entries that contain "Champion" or "1st Round" (if applicable).
Step 3: SQL - Filter the data to find the first year after 1936 where the National_Cup column
indicates a win. Ensure to handle any NaN or ambiguous entries appropriately.
@@ A.4 Code Execution Results
Listing 13: Response to Sample Prompt 4 (Code Execution Prompt) Step 1
@@
- LLM prompt: Extract the year from phrases like "Spring 1932" or "Fall 1932" and standardize all
entries to a YYYY format. Ensure the output is consistent across all entries.
@@
Listing 14: Response to Sample Prompt 4 (Code Execution Prompt) Step 2
@@
dataframe create after loading from standardized_national_cup ...
@@
Listing 15: Response to Sample Prompt 4 (Code Execution Prompt) Step 3
@@
dataframe create after loading from first_win_after_1936 ...
@@
Listing 16: Response to Sample Prompt 5 (Answer Extraction Prompt)
@@ A.5 More Detail on Optimization
Listing 17: Example of Step Merging
@@
Parallel LLM Execution.
Initially, we prompted the LLM to generate a new column by supplying the entire existing column and asking it to return a list of the same length. However, this approach often led to inconsistent results—such as incorrect list lengths, duplicated values, or hallucinated entries—due to the model’s sensitivity to long input sequences. Errors were especially prevalent in the middle of the list, consistent with the “Lost in the Middle” effect Liu et al. (2023).
@@
To improve both reliability and efficiency, we adopted a chunk-wise parallel execution strategy that avoids the overhead of row-by-row processing while enhancing consistency. As illustrated in Figure 5, we segment the input into appropriately sized batches and execute multiple LLM calls in parallel. This design enables simultaneous inference over different parts of the data, substantially reducing latency by eliminating sequential processing bottlenecks. The result is faster response times and improved scalability—making this approach well-suited for large-scale reasoning tasks over semi-structured data.
Column-Wise Batching. Conventional LLM pipelines often chunk inputs row-wise, generating one column value per row across a batch. In contrast, we propose a column-wise batching strategy, depicted in Figure 6, which processes multiple columns for a small chunk of rows in a single call.
@@
Listing 18: Examples of Different Plan Optimization
Question - The Kremlin Cup is held in Russia, and the St. Petersburg Open is also held in Russia.
@@
Step 1: SQL - Filter the table to select tournaments with the names "Kremlin Cup" and "St. Petersburg Open".
Step 2: SQL - Extract the country information from the Tournament column for the selected tournaments.
Step 3: LLM - Summarize the results to confirm that both tournaments are held in Russia.
@@
Step 1: SQL - Filter the table to select tournaments with the names "Kremlin Cup" and "St. Petersburg Open", and extract the country information from the Tournament column in a single query.
Step 2: LLM - Summarize the results to confirm that both tournaments are held in Russia.
A.6 More Detail on handling Multi-Modal data
The proposed pipeline Figure 8 is modular and designed for extensibility. Each component can be upgraded such as substituting the SQL Query Executor with an expert SQL agent to enhance execution efficiency and accuracy. Likewise, the LLM Semantic Reasoner and VLM (Vision Language Model) components can be replaced with specialized reasoning agents, allowing Weaver to adapt to evolving multi-modal requirements.
@@ A.6 More Detail on handling Multi-Modal data
Handling Paragraph data using LLM step
@@
Figure 8: Modular Pipeline for query execution
@@ A.7 Additional Analysis
Token Usage/Query GPT-4o (I/O) GPT-4o-mini (I/O) Gemini-2.0 (I/O)

### v2

@@ 3.2 Results and Analysis
Table 4: Number of API calls comparison per Table QA. : Gemini-2.0 refers to Gemini-2.0-Flash.
ProTrix uses only two fixed API calls and relies on the LLM solely for generating the plan. However, it does not involve the LLM during execution. This limits its ability to handle multi-step queries requiring reasoning at each step. For instance, it may fail to infer information from individual rows or perform operations such as applying a SQL GROUP BY on an LLM-inferred country column. Such steps are often essential to arrive at the correct final answer. These metrics demonstrate how our approach minimizes computational overhead while maintaining accuracy.
Optimization:
We experimented with optimizing our planning and execution strategy on 200 Table QA queries, which were randomly sampled from the evaluation benchmarks. Table 5 demonstrates how our optimization strategy focuses on reducing unnecessary computational steps without compromising accuracy.
@@
This optimization was achieved by targeting redundancies in both the LLM and SQL steps.
1. LLM Step Reduction:
By identifying and eliminating redundant LLM steps, we reduced the 15 LLM calls. This optimization ensures that LLMs are only used when necessary, lowering computational costs.
2. SQL Step Optimization:
We achieved a reduction of 19 SQL steps by eliminating unused operations. Additionally, we merged 113 sequential SQL steps into fewer, more efficient queries and reordered 4 steps to optimize execution flow, making it more efficient.
@@ 2. SQL Step Optimization:
Table 6: Error in SQL and Plan Generation on WikiTQ. : Gemini-2.0 refers to Gemini-2.0-Flash.
Error Analysis:
@@
We perform a stage-wise analysis to assess the contribution of each component in our pipeline: filtering, planning, and execution. Compared to SQL-only generation, which struggles with multi-step reasoning, our pipeline yields consistent accuracy gains by structuring the task into sub-components. The filtering stage removes irrelevant columns, reducing noise and guiding the model’s attention. The planning stage, central to our method, decomposes complex queries into symbolic and semantic steps. This step is essential and cannot be ablated. However, skipping plan verification (i.e., executing without validating) leads to a 1% drop in accuracy. Finally, execution stage translates structured plans into SQL query and LLM prompts with column details.
4 Comparison with Related Work
Table-based Question Answering (Table QA) combines table understanding, question interpretation, and NLP. Foundational work such as Text-to-SQL Tan et al. (2024), Program-of-Thought Chen et al. (2023), and TabSQLify Nahid and Rafiei (2024b) laid the groundwork. Binder and TAG Biswal et al. (2024) expose the limitations of traditional Text-to-SQL methods in handling complex analytical tasks involving both structured and unstructured data. To address these challenges, several alternative approaches have been explored:
Fine-Tuning Methods:
These methods fine-tune LLMs to specialize in reasoning over hybrid tabular and textual data. Models such as Zhu et al. (2024), Mittal et al. (2024), and Patel et al. (2024) are trained to extract, reason, and execute over such inputs. However, fine-tuning requires large task-specific datasets and tends to lack generalization across domains.
Query Engines:
This direction integrates LLM with SQL engines via user-defined functions (UDFs), allowing LLM calls within queries. UQE Dai et al. (2024), BlendSQL, SUQL Liu et al. (2024b), and Binder follow this paradigm. While flexible, LLM-generated queries can be error-prone, and these systems often support only limited query structures, reducing adaptability.
Table Pruning and Planning:
Approaches like H-STAR, ReAcTable, ProTrix, and others Liu et al. (2024c); Nahid and Rafiei (2024a) enhance efficiency by programmatically pruning rows or columns using SQL or Python. While this reduces processing overhead, these methods often function as black boxes, lacking transparency and vulnerable to cascading errors if early pruning steps are incorrect. More information in Appendix F.
@@ 5 Conclusion
We introduce Weaver , a novel approach for table-based question answering on tables with embedded unstructured text. Weaver outperforms all baselines by strategically decomposing complex queries into a sequence of LLM- and SQL-based planning steps. By alternating between these modalities, it enables precise, interpretable, and adaptive query resolution. Weaver overcomes prior limitations by effectively handling both complex queries and large tables. Its modular design also supports future extensions, including image-based tables, multi-table reasoning, and integration with free-form text. As future work, we plan to explore fine-tuning and supervision strategies to further improve execution accuracy and plan reliability.
@@ Acknowledgments
We gratefully acknowledge the Cognitive Computation Group at the University of Pennsylvania and the Complex Data Analysis and Reasoning Lab at Arizona State University for their resources and computational support. A part of this work was funded by ONR Contract N00014-23-1-2364 and N00014-23-1-2417. Yanjie is supported by the National Science Foundation (NSF) via the grant numbers: 2426340, 2416727, 2421864, 2421865, 2421803, and National academy of engineering Grainger Foundation Frontiers of Engineering Grants. We also thanks the reviewers for there thoughtful feedback and comment.
@@ 2. The list will be used to filter the table dataframe directly so take care of that
- Check if the MySQL step logic adheres to the column format. (Performs calculations
and formatting and filtering in the table)
@@
Do not provide code or other explanations, only the new plan.
@@
1. Store the output at each step by creating a new table. Use this new table for the
next steps.
2. The code for MySQL should handle all values in the column (formatting and
filtering). New columns
@@ 2. The code for MySQL should handle all values in the column (formatting and
LLM step format should be the same.
@@
Give me code for solving the question, and no other explanations. Keep in mind the
column data formats (string to appropriate data type, removing extra character, Null
values) while writing Mysql code.
@@
- Do not provide any explanation and Return only a list (separate values by ’#’)
that can be added
@@
- Any value should not be more than 3 words (or each value should be as short as
possible).
@@
If the table has just one column or value consider that as the answer given the
column name.
@@
Listing 8: Answer Format Prompt
Table: {self.name}
{self.table}
Question: {self.question}
You will be given Answer and Gold Answer, you have to Convert the answer into a
format of gold answer given above, if the content or meaning is same (semantically
same) they should be same.
Few examples of conversion for your understanding:
1. answer: ITA, gold answer: Italy. Reasoning- ITA is country code of Italy hence ITA and Italy are same and you can convert ITA to Italy.
Your Output: Italy
2. answer: 17, gold answer: 17 years. Reasoning- 17 of answer is same as 17 years of the gold answer in the given context of question.
Your Output: 17 years
3. answer : 10, gold answer: 10. Reasoning- Since, both values are already same no conversion is needed.
Your Output: 10
4. answer : 0, gold answer: 5. Reasoning- Since, both values are semantically not same no conversion is needed for the answer.
Your Output: 0
5. answer : The answer is not present in the table. , gold answer: 5. Reasoning- Since, both values are semantically not same no convertion is needed for the answer.
Your Output: The answer is not present in the table.
Listing 9: Plan Optimization Prompt
@@
SQL Merge: Merge sequential SQL steps where possible (e.g., combining filtering,
aggregation, and sorting in one query).
SQL Reordering: Reorder SQL steps to filter early before applying computationally
expensive operations like LLM processing.
@@ A.2 Table and Question Example
Listing 10: Table and Question Example for LLM Planning
@@
Question: How long did it take for the New York Americans to win the National Cup
after 1936?
@@ A.3 Model Responses
Listing 11: Response to Sample Prompt 1 (Column Description)
@@
The table New_York_Americans_soccer contains historical performance data for the
New York Americans soccer team, detailing their standings in various leagues,
playoff outcomes, and national cup results from 1931 to 1956. The data includes
information on the year, division, league, regular season standings, playoff
participation, and national cup results.
@@
| Year | String | Standardize to a consistent format
| Represents the year or season of the soccer performance. Some entries may need parsing to extract the year. |
| Division | Float | Convert to Integer (if applicable)
| Indicates the division in which the team played. Some entries are NaN and need handling. |
| League | String | Standardize to a consistent format
| Represents the league in which the team participated, primarily ASL (American Soccer League). |
| Reg_Season | String | Clean and standardize standings
| Indicates the teams regular season performance, which may include qualifiers like "t" or "?" that need clarification. |
| Playoffs | String | Standardize and clean
| Indicates playoff participation and outcomes, with entries like "No playoff" and "Finals" that need to be standardized. |
| National_Cup | String | Standardize and clean
| Indicates the outcome of the national cup, with entries like "Champion" or "1st Round" that need standardization. |
Listing 12: Response to Sample Prompt 2 (Planning Prompt)
Plan: Step 1: SQL - Standardize the Year column to a consistent format and extract
the year from entries like "Spring 1932" and "Fall 1932".
Step 2: SQL - Clean and standardize the National_Cup column to identify the years
when the team won the national cup.
Step 3: SQL - Filter the data to find the first year after 1936 when the
National_Cup column indicates a win.
Listing 13: Response to Sample Prompt 3 (Verify Plan Prompt)
@@
Step 1: LLM - Standardize the Year column to a consistent format by extracting the
year from entries like "Spring 1932" and "Fall 1932". Convert all entries to a four-
digit year format (e.g., "1932" instead of "Spring 1932").
Step 2: SQL - Clean and standardize the National_Cup column to identify winning
entries. Define a clear criterion for a "win," such as entries that contain
"Champion" or "1st Round" (if applicable).
Step 3: SQL - Filter the data to find the first year after 1936 where the
National_Cup column indicates a win. Ensure to handle any NaN or ambiguous entries
appropriately.
@@ A.4 Code Execution Results
Listing 14: Response to Sample Prompt 4 (Code Execution Prompt) Step 1
@@
- LLM prompt: Extract the year from phrases like "Spring 1932" or "Fall 1932" and
standardize all entries to a YYYY format. Ensure the output is consistent across all
entries.
@@
Listing 15: Response to Sample Prompt 4 (Code Execution Prompt) Step 2
@@
Dataframe created after loading from standardized_national_cup ...
@@
Listing 16: Response to Sample Prompt 4 (Code Execution Prompt) Step 3
@@
Dataframe created after loading from first_win_after_1936 ...
@@
Listing 17: Response to Sample Prompt 5 (Answer Extraction Prompt)
@@ A.5 More Detail on Optimization
Listing 18: Example of Step Merging
@@
Parallel LLM Execution:
Initially, we prompted the LLM to generate a new column by supplying the entire existing column and asking it to return a list of the same length. However, this approach often led to inconsistent results, such as incorrect list lengths, duplicated values, or hallucinated entries, due to the model’s sensitivity to long input sequences. Errors were especially prevalent in the middle of the list, consistent with the “Lost in the Middle” effect Liu et al. (2024a).
@@
To improve both reliability and efficiency, we adopted a chunk-wise parallel execution strategy that avoids the overhead of row-by-row processing while enhancing consistency. As illustrated in Figure 5, we segment the input into appropriately sized batches and execute multiple LLM calls in parallel. This design enables simultaneous inference over different parts of the data, substantially reducing latency by eliminating sequential processing bottlenecks. The result is faster response times and improved scalability, making this approach well-suited for large-scale reasoning tasks over semi-structured data.
Column-Wise Batching: Conventional LLM pipelines often chunk inputs row-wise, generating one column value per row across a batch. In contrast, we propose a column-wise batching strategy, depicted in Figure 6, which processes multiple columns for a small chunk of rows in a single call.
@@
Weaver implements this with a fixed maximum batch size defined in terms of the total number of values (columns × rows). To handle cases where the number of selected columns is large and could exceed the context length, the strategy adaptively adjusts the number of rows based on the number of columns (e.g., if more columns are included, fewer rows are processed per batch, while a smaller number of columns allows more rows). This adaptive trade-off ensures that the total input length remains manageable and consistent across diverse table structures.
@@
Listing 19: Examples of Different Plan Optimization
Question - The Kremlin Cup is held in Russia, and the St. Petersburg Open is also
held in Russia.
@@
Step 1: SQL - Filter the table to select tournaments with the names "Kremlin Cup"
and "St. Petersburg Open".
Step 2: SQL - Extract the country information from the Tournament column for the
selected tournaments.
Step 3: LLM - Summarize the results to confirm that both tournaments are held in
Russia.
@@
Step 1: SQL - Filter the table to select tournaments with the names "Kremlin Cup"
and "St. Petersburg Open", and extract the country information from the Tournament
column in a single query.
Step 2: LLM - Summarize the results to confirm that both tournaments are held in
Russia.
A.6 More Detail on handling Multimodal data
The proposed pipeline Figure 7 is modular and designed for extensibility. Each component can be upgraded such as substituting the SQL Query Executor with an expert SQL agent to enhance execution efficiency and accuracy. Likewise, the LLM Semantic Reasoner and VLM (Vision Language Model) components can be replaced with specialized reasoning agents, allowing Weaver to adapt to evolving multimodal requirements.
@@ A.6 More Detail on handling Multimodal data
Processing Paragraph:
@@
Figure 8: Paragraph-aware Weaver pipeline
@@ A.7 Additional Analysis
Token Usage/Query GPT-4o (I/O) GPT-4o-mini (I/O) Gemini-2.0-Flash (I/O)